# A Survey of Fuzzing Open-Source Operating Systems

### v2

@@ 3.3. OSF Workflow
We summarize the general workflow of OSF in Figure 4, which consists of three key modules, i.e., input, fuzzing engine, and running environment. Each module is described in detail as follows.
Figure 4. General Workflow of OSF (It has three modules, i.e., input, fuzzing engine, and running environment).
@@
Input. There are three types of inputs to be considered in the input module, i.e., initial seed, specification, and target OS. First, the initial seed is the raw material fed to the fuzzing engine, which is initially deposited in a global corpus (i.e., a repository for storing seed candidates) and subsequently used throughout the fuzzing workflow. Second, the specification, which describes the structure and syntax of the seed, is parsed by the fuzzing engine and used to generate new seeds or enhance the quality of the seeds. It is important to note that the use of the specifications is not applicable to all scenarios; it mainly applies to highly structured seeds with documented descriptions of the seed structure (Google, 2015; Chen et al., 2020b; Zhao et al., 2022a; Sun et al., 2021; Shen et al., 2021, 2022b; Jeong et al., 2023b; Xu et al., 2020; Jeong et al., 2019). Third, the target OS requires to be instrumented to collect runtime information for feedback analysis and bug monitoring during fuzzing.
@@
Fuzzing Engine. The main function of the fuzzing engine is to generate new seeds and wrap them into test cases for the input fed to the executor in the running environment. This module go through a complete closed loop of seed selection from a corpus, seed mutation and/or generation, and seed trim before storing to the corpus. Specifically, the corpus stores only seeds that have been verified as “high-quality”, where “high-quality” seeds can be defined as the seeds that have triggered bugs (Zou et al., 2022; Wu et al., 2018; Lin et al., 2022; Shen et al., 2021), the initial seed generated based on tailored rules (Shen et al., 2022a) or specifications (Google, 2015), or the streamlined seeds after seed trim as they trigger new coverage (Pailoor et al., 2018; Shen et al., 2021). Such a consideration stems from the empirical conclusion that mutations based on high-quality seeds usually have more opportunities to move the execution of the PUT closer to trigger bugs (You et al., 2017; Shen et al., 2022a; Lin et al., 2022; Maier et al., 2019). However, even if the corpus always stores high-quality seeds, it is necessary to consider which seeds should be prioritized in order to improve the efficiency of seed mutation. Therefore, seed selection is often adopted to prioritize the seeds in the corpus, and a computation method of seed priority is often devised to ensure that the selected seeds have more chances to reveal new bugs, e.g., PageRank (Lin et al., 2022), program analysis (Wu et al., 2018; Chen et al., 2022a), evolutionary algorithm (Shen et al., 2022a; Henderson et al., 2017; Xu et al., 2020), empirical strategy (Xu et al., 2020; Google, 2015; You et al., 2017; Zou et al., 2022), reinforcement learning (Wang et al., 2021), etc. Then, to continuously generate test cases to automate the whole fuzzing process, seed mutation and/or generation transform and/or generate the seeds in some way and wrap them into test cases that can be executed by the target OS. For example, seed mutation employs parameter-level bit flips or byte-level replacement to generate seeds based on the high-quality seeds in the corpus. Seed generation builds test case models from the specification to generate new seeds. Finally, to further improve the speed performance of OSF, the generated seed usually requires to be trimmed to reduce its mutation space as well as its execution time, e.g., greedy algorithm (Pailoor et al., 2018) and stepwise filtering (Schumilo et al., 2020; Henderson et al., 2017). The trimmed seeds are then stored into the corpus, and the process is repeated to automatically drive new and high-quality seed generation.
@@
Running Environment. The running environment is responsible for feeding the test cases wrapped by the fuzzing engine into the executor. Unlike traditional fuzzing techniques, OSF usually relies on full-virtual environment to avoid as much as possible the effects propagated by OS crashes. In contrast, traditional fuzzing techniques do not fatally affect the entire fuzzing process even if a crash occurs. Therefore, the running environment of OSF is often designed as full-virtual environment that can be quickly recovered, allowing the entire system to be quickly restarted to a known, clean state in the event of a crash (Shen et al., 2022b; Lin et al., 2022; Schumilo et al., 2021; Pan et al., 2021; Zhao et al., 2022b; Song et al., 2020). Such a design allows a fuzzer to run thousands of test cases continuously and automatically in an isolated environment without worrying about the potential for lasting damage from individual test cases. Subsequent execution results fall into two categories. If the executor does not crash, the feedback will be collected by the instrumented code, which is used to guide seed mutation of the fuzzing engine to steer the fuzzing towards uncovered program paths. If the executor triggers a crash, the instrumented code records the information and sends it to the bug analyzer. The bug analyzer identifies and categorizes errors, and generates detailed bug reports. Finally, if the fuzzing process is not yet finished, it should be restored to its pre-crash state, and then move on to explore other states of the target OS.
@@
Although there are some similarities with traditional fuzzing in terms of the coarse-grained fuzzing process, the technical challenges and implementation details of each step need to take into account the specificity and complexity of OS, which ultimately makes it show a big difference with traditional fuzzing in terms of the fine-grained details. Therefore, we will elaborate on the technical details and technical distinctions of each step illustrated in Figure 4 in Section 4 and 5.
4. OSF Steps In Detail
OSF differs significantly from traditional fuzzing in technical approach, largely due to the OS’s intricate modular design, its extensive concurrency, and the complex, diverse interaction interfaces connecting user space and hardware with kernel space. These structural complexities create unique challenges for fuzzing, impacting both its effectiveness and efficiency in practical implementations. Thus, we review the details of the three core modules—input, fuzzing engine, and runtime environment—shown in Figure 4, and comprehensively compare the applicability scenarios of various fuzzing techniques.
Table 1. Overview of Fuzzers Sorted by Publication Year.
Fuzzer Input Fuzzing Engine Running Environment
OSL ISG I. SS ST U. F. VTs
Trinity(D.Jones, 2015) Kernel Spec. - - - Gen.(1) Code -
Syzkaller(Google, 2015) Kernel Spec.’ Static Feedback Feedback Mut. Code -
KernelFuzzer(MWR Labs, 2016) Kernel Spec. Static - - Gen.(1) Code -
DIFUZE(Corina et al., 2017) Driver Spec.’ Static - - Gen.(2) - M.,L.,O.
VDF(Henderson et al., 2017) Hypervisor Trace Static - Crash Mut. Code M.,C.,L.
KAFL(Schumilo et al., 2017) File System - Static+Dynamic - - Mut. Code M.
Semfuzz(You et al., 2017) Kernel PoC Static Distance - Mut. Code M.
usb-fuzzer(Google, 2015) Driver Spec.’ Static - - Mut. Code M.
Moonshine(Pailoor et al., 2018) Kernel Trace Static Feedback - Mut. Code M.,C.
FUZE(Wu et al., 2018) Kernel PoC Static - Distance Mut. Code M.
Schwarz et al.(Schwarz et al., 2018) Kernel Trace - - - Gen.(1) Code C.
Razzer(Jeong et al., 2019) Kernel Spec.’ Static - - Mut. Code C.
JANUS(Xu et al., 2019) File System - Static - - Mut. Code+Cust. M.,L.
SLAKE(Chen and Xing, 2019) Kernel PoC Static - Distance Mut. - M.
Shi et al.(Shi et al., 2019) Kernel Spec.’ Static - - Mut. - M.,C.,O.
PeriScope(Song et al., 2019a) Driver Trace Static - Feedback Mut. Code M.
Unicorefuzz(Maier et al., 2019) Kernel Pattern Static - - Mut. Code M.
KOOBE(Chen et al., 2020b) Kernel PoC Dynamic - - Mut. Code+Cust. M.
@@ 4. OSF Steps In Detail
Krace(Xu et al., 2020) File System Spec.’ Static Minimum - Mut. Code+Thread C.
Hyper-CUBE(Schumilo et al., 2020) Hypervisor Pattern - - Crash - - M.,C.,L.,O.
Hydra(Kim et al., 2020b) File System - Static - Crash Mut. Code+Cust. M.,L.
USBFuzz(Peng and Payer, 2020) Driver Pattern Static - - Mut. Code M.
Agamotto(Song et al., 2020) Driver Trace Static - - Mut. Code -
Ex-vivo(Pustogarov et al., 2020) Driver Trace Static - - Mut. Code M.
HFL(Kim et al., 2020a) Kernel Spec.’ Static - - Mut. Code M.
X-AFL(Liang et al., 2020) Kernel Trace Static - - Mut. Code M.
HEALER(Sun et al., 2021) Kernel Spec.’ Static - Feedback Gen.+Mut. Code M.,C.,L.
Rtkaller(Shen et al., 2021) Kernel Spec.’ Static - - Mut. Code M.,C.
NYX(Schumilo et al., 2021) Hypervisor Spec. Dynamic - - Mut. Code M.
V-Shuttle(Pan et al., 2021) Hypervisor Trace Static - - Mut. Code M.,L.
BSOD(Maier and Toepfer, 2021) Driver Pattern Dynamic - - Gen.+Mut. Code M.
SyzVegas(Wang et al., 2021) Kernel Spec.’ Static Feedback - Mut. Code M.
@@
StateFuzz(Zhao et al., 2022a) Driver Spec.’ Static Similarity - Mut. Code+Cust. M.
GREBE(Lin et al., 2022) Kernel PoC Static Minimum - Mut. Custom M.,O.
MundoFuzz(Myung et al., 2022) Hypervisor Trace Static - - Mut. Code M.,L.
Morphuzz(Bulekov et al., 2022) Hypervisor Trace Static - - Mut. Code C.,L.
@@
CONZZER(Jiang et al., 2022) File System Pattern Static Feedback - Mut. Code+Thread C.
@@
Dr.Fuzz(Zhao et al., 2022b) Driver Spec.’ Static+Dynamic - - Mut. Code+Cust. M.,O.
PrintFuzz(Ma et al., 2022) Driver Spec.’ Static+Dynamic - - Gen.+Mut. Code M.,C.
DriFuzz(Shen et al., 2022a) Driver Trace Static - - Gen.+Mut. Code M.
Hao et al.(Hao et al., 2022) Kernel - Static - - - - -
KSG(Sun et al., 2022) Kernel Spec.’ Static - - - - M.,C.,L.
SyzScope(Zou et al., 2022) Kernel PoC Static Feedback - Mut. - M.
Tardis(Shen et al., 2022b) Kernel Spec. Static - - Mut. Code M.,L.
Segfuzz(Jeong et al., 2023b) Kernel Spec.’ Static - - Mut. Code+Thread M.,C.,O.
IRIS(Cesarano et al., 2023) Hypervisor Trace Static - - Mut. Code M.
FUZZNG(Bulekov et al., 2023) Kernel Pattern Static - - Mut. Code M.
ACTOR(Fleischer et al., 2023) Kernel Trace Static - - Gen.(2) Code M.,L.
SyzDescribe(Hao et al., 2023) Driver Spec.’ - - - - - -
DEVFUZZ(Wu et al., 2023) Driver Pattern Static+Dynamic - - Mut. Code+Cust. M.,L.,O.
Syzdirect(Tan et al., 2023) Kernel Spec.’ Static Distance Distance Mut. Code -
ReUSB(Jang et al., 2023) Driver Trace Dynamic - - Mut. Code M.

### v3

@@ 3.3. OSF Workflow
We summarize the general workflow of OSF in Figure 3(b), which consists of three key modules, i.e., input, fuzzing engine, and running environment. Each module is described in detail as follows.
@@
Input. There are three types of inputs to be considered in the input module, i.e., initial seed, specification, and target OS. First, the initial seed is the raw material fed to the fuzzing engine, which is initially deposited in a global corpus (i.e., a repository for storing seed candidates) and subsequently used throughout the fuzzing workflow. Second, the specification, which describes the structure and syntax of the seed, is parsed by the fuzzing engine and used to generate new seeds or enhance the quality of the seeds. It is important to note that the use of the specifications is not applicable to all scenarios; it mainly applies to highly structured seeds with documented descriptions of the seed structure (Syzkaller; chen2020koobe; zhao2022statefuzz; sun2021healer; shen2021rtkaller; shen2022tardis; jeong2023segfuzz; xu2020krace; jeong2019razzer). Third, the target OS requires to be instrumented to collect runtime information for feedback analysis and bug monitoring during fuzzing.
@@
Fuzzing Engine. The main function of the fuzzing engine is to generate new seeds and wrap them into test cases for the input fed to the executor in the running environment. This module go through a complete closed loop of seed selection from a corpus, seed mutation and/or generation, and seed trim before storing to the corpus. Specifically, the corpus stores only seeds that have been verified as “high-quality”, where “high-quality” seeds can be defined as the seeds that have triggered bugs (zou2022syzscope; wu2018fuze; lin2022grebe; shen2021rtkaller), the initial seed generated based on tailored rules (shen2022drifuzz) or specifications (Syzkaller), or the streamlined seeds after seed trim as they trigger new coverage (pailoor2018moonshine; shen2021rtkaller). Such a consideration stems from the empirical conclusion that mutations based on high-quality seeds usually have more opportunities to move the execution of the PUT closer to trigger bugs (you2017semfuzz; shen2022drifuzz; lin2022grebe; Unicorefuzz2019). However, even if the corpus always stores high-quality seeds, it is necessary to consider which seeds should be prioritized in order to improve the efficiency of seed mutation. Therefore, seed selection is often adopted to prioritize the seeds in the corpus, and a computation method of seed priority is often devised to ensure that the selected seeds have more chances to reveal new bugs, e.g., PageRank (lin2022grebe), program analysis (wu2018fuze; chen2022sfuzz), evolutionary algorithm (shen2022drifuzz; henderson2017vdf; xu2020krace), empirical strategy (xu2020krace; Syzkaller; you2017semfuzz; zou2022syzscope), reinforcement learning (wang2021syzvegas), etc. Then, to continuously generate test cases to automate the whole fuzzing process, seed mutation and/or generation transform and/or generate the seeds in some way and wrap them into test cases that can be executed by the target OS. For example, seed mutation employs parameter-level bit flips or byte-level replacement to generate seeds based on the high-quality seeds in the corpus. Seed generation builds test case models from the specification to generate new seeds. Finally, to further improve the speed performance of OSF, the generated seed usually requires to be trimmed to reduce its mutation space as well as its execution time, e.g., greedy algorithm (pailoor2018moonshine) and stepwise filtering (schumilo2020hyper; henderson2017vdf). The trimmed seeds are then stored into the corpus, and the process is repeated to automatically drive new and high-quality seed generation.
@@
Running Environment. The running environment is responsible for feeding the test cases wrapped by the fuzzing engine into the executor. Unlike traditional fuzzing techniques, OSF usually relies on full-virtual environment to avoid as much as possible the effects propagated by OS crashes. In contrast, traditional fuzzing techniques do not fatally affect the entire fuzzing process even if a crash occurs. Therefore, the running environment of OSF is often designed as full-virtual environment that can be quickly recovered, allowing the entire system to be quickly restarted to a known, clean state in the event of a crash (shen2022tardis; lin2022grebe; schumilo2021nyx; pan2021V-shuttle; zhao2022semantic; song2020agamotto). Such a design allows a fuzzer to run thousands of test cases continuously and automatically in an isolated environment without worrying about the potential for lasting damage from individual test cases. Subsequent execution results fall into two categories. If the executor does not crash, the feedback will be collected by the instrumented code, which is used to guide seed mutation of the fuzzing engine to steer the fuzzing towards uncovered program paths. If the executor triggers a crash, the instrumented code records the information and sends it to the bug analyzer. The bug analyzer identifies and categorizes errors, and generates detailed bug reports. Finally, if the fuzzing process is not yet finished, it should be restored to its pre-crash state, and then move on to explore other states of the target OS.
@@
Although there are some similarities with traditional fuzzing in terms of the coarse-grained fuzzing process, the technical challenges and implementation details of each step need to take into account the specificity and complexity of OS, which ultimately makes it show a big difference with traditional fuzzing in terms of the fine-grained details. Therefore, we will elaborate on the technical details and technical distinctions of each step illustrated in Figure 3(b) in Section 4 and LABEL:Section5.
4. General Workflow Characteristics of OSF
We review the state-of-the-art OSF with respect to the key steps in the three core modules of the general fuzzing workflow, i.e., input, fuzzing engine, and runtime environment, as shown in Figure 3(b). Our review results are reported in Table LABEL:tab1, where “-” means that this dimension is not relevant, or not clearly mentioned in the fuzzer; ISG denotes initial seed generation, Inst. denotes instrumentation, SS, ST and SU respectively denote seed selection, seed trim, and seed update (generation or mutation); Feed. denotes feedback mechanism; and BT denotes supported bug types.
Table 1. Characteristics of OS Fuzzers (Sorted by Publication Year)
Trinity(Trinity) Kernel Spec. - - - Gen. Code -
Syzkaller(Syzkaller) Kernel Spec.’ Static Feedback $D$ Mut. Code -
TriforceAFL(TriforceAFL) Kernel Pattern Dynamic Feedback $D$ Mut. Code M.
DIFUZE(2017DIFUZE) Driver Spec.’ Static - - Gen. - M.,L.,O.
VDF(henderson2017vdf) Hypervisor Trace Static - $D$ Mut. Code M.,C.,L.
kAFL(schumilo2017kafl) File System - Static+Dynamic - - Mut. Code M.
SemFuzz(you2017semfuzz) Kernel PoC Static Distance - Mut. Code M.
usb-fuzzer(Syzkaller) Driver Spec.’ Static - - Mut. Code M.
Moonshine(pailoor2018moonshine) Kernel Trace Static Feedback - Mut. Code M.,C.
FUZE(wu2018fuze) Kernel PoC Static - - Mut. Code M.
Schwarz et al.(schwarz2018automated) Kernel Trace - - - Gen. Code C.
Razzer(jeong2019razzer) Kernel Spec.’ Static - - Mut. Code C.
JANUS(JANUS2019fuzzing) File System - Static - - Mut. Code+Cust. M.,L.
SLAKE(chen2019slake) Kernel PoC Static - - Mut. - M.
Shi et al.(shi2019industry) Kernel Spec.’ Static - - Mut. - M.,C.,O.
PeriScope(Song2019PeriScopeAE) Driver Trace Static - - Mut. Code M.
Unicorefuzz(Unicorefuzz2019) Kernel Pattern Static - - Mut. Code M.
KOOBE(chen2020koobe) Kernel PoC Dynamic - - Mut. Code+Cust. M.
@@ 4. General Workflow Characteristics of OSF
Krace(xu2020krace) File System Spec.’ Static Minimum - Mut. Code+Thread C.
Hyper-Cube(schumilo2020hyper) Hypervisor Pattern - - $D$ - - M.,C.,L.,O.
Hydra(Hydra2020finding) File System - Static - - Mut. Code+Cust. M.,L.
USBFuzz(peng2020usbfuzz) Driver Pattern Static - - Mut. Code M.
Agamotto(song2020agamotto) Driver Trace Static - - Mut. Code -
Ex-vivo(Pustogarov2020ExvivoDA) Driver Trace Static - - Mut. Code M.
HFL(kim2020hfl) Kernel Spec.’ Static - - Mut. Code M.
X-AFL(liang2020xafl) Kernel Trace Static - - Mut. Code M.
HEALER(sun2021healer) Kernel Spec.’ Static - - Gen.+Mut. Code M.,C.,L.
Rtkaller(shen2021rtkaller) Kernel Spec.’ Static - - Mut. Code M.,C.
NYX(schumilo2021nyx) Hypervisor Pattern Dynamic - - Mut. Code M.
V-Shuttle(pan2021V-shuttle) Hypervisor Trace Static - - Mut. Code M.,L.
BSOD(maier2021bsod) Driver Pattern Dynamic - - Mut. Code M.
SyzVegas(wang2021syzvegas) Kernel Spec.’ Static Feedback - Mut. Code M.
@@
StateFuzz(zhao2022statefuzz) Driver Spec.’ Static Similarity - Mut. Code+Cust. M.
GREBE(lin2022grebe) Kernel PoC Static Minimum - Mut. Cust. M.,O.
MundoFuzz(myung2022mundofuzz) Hypervisor Trace Static - - Mut. Code M.,L.
Morphuzz(bulekov2022morphuzz) Hypervisor Trace Static - - Mut. Code C.,L.
@@
CONZZER(CONZZER2022context) File System Pattern Static Feedback - Mut. Code+Thread C.
@@
Dr.Fuzz(zhao2022semantic) Driver Spec.’ Static+Dynamic - - Mut. Code+Cust. M.,O.
PrIntFuzz(ma2022printfuzz) Driver Spec.’ Static+Dynamic - - Gen.+Mut. Code M.,C.
DriFuzz(shen2022drifuzz) Driver Trace Static - - Gen.+Mut. Code M.
Hao et al.(Hao2022DemystifyingTD) Kernel - Static - - - - -
KSG(sun2022ksg) Kernel Spec.’ Static - - - - M.,C.,L.
SyzScope(zou2022syzscope) Kernel PoC Static - - Mut. - M.
Tardis(shen2022tardis) Kernel Spec. Static - - Mut. Code M.,L.
SegFuzz(jeong2023segfuzz) Kernel Spec.’ Static - - Mut. Code+Thread M.,C.,O.
IRIS(cesarano2023iris) Hypervisor Trace Static - - Mut. Code M.
FUZZNG(bulekov2023FUZZNG) Kernel Pattern Static - - Mut. Code M.
ACTOR(fleischer2023actor) Kernel Trace Static - - Gen. Code+Cust. M.,L.
SyzDescribe(hao2023syzdescribe) Driver Spec.’ - - - - - -
DEVFUZZ(wu2023devfuzz) Driver Pattern Static+Dynamic - - Mut. Code+Cust. M.,L.,O.
SyzDirect(tan2023syzdirect) Kernel Spec.’ Static Distance - Mut. Code -
ReUSB(Jang2023ReUSB) Driver Trace Dynamic - Protocol Mut. Code M.